I'm a 7th grader in pre-Algebra, and I need help with a multi-step equation. The teacher told us to use inverse operations and inverse order of operations to solve this equation:
y + 9
_____ = 2.1
4
Or (y+9)/4=2.1 And I'm hoping to find Y and how you solve this equation using
inverse operations. Thank you! -- Josh ---

Answer by richwmiller(17219)

the inverse of 1/4 is 4
multiply both sides by 4
y+9=2.1*4
y+9=8.4
add the inverse of 9 (-9)
y=-.6